The lucida package^{∗}

Sebastian Rahtz s.rahtz@elsevier.co.uk 1994/10/02

∗This file has version number v1.5i, last revised 1994/10/02.

March 12, 1996

1 Introduction

This file contains L^{A}TEX 2ε package files needed to use Lucida Bright fonts, and .fd files for the fonts as named by Y&Y. .fd and metric files for the Berry-naming scheme are available in fonts/metrics/bh on CTAN.

The Lucida Bright font families:

hlxd8a  lfd  LucidaFax-Demi
hlxdi8a lfdi  LucidaFax-DemiItalic
hlxr8a lfr LucidaFax
hlxri8a lfi  LucidaFax-Italic
hlhd8a  lbd  LucidaBright-Demi
hlhdi8a lbdi  LucidaBright-DemiItalic
hlhr8a lbr LucidaBright
hlhri8a lbi  LucidaBright-Italic
hlhro8a lbsl  LucidaBrightSlanted
hlhrc8a lbrsc LucidaBrightSmallcaps
hlhdc8a lbdsc LucidaBrightSmallcaps-Demi
hlsdi8a lsdi  LucidaSans-DemiItalic
hlsd8a  lsd  LucidaSans-Demi
hlsri8a  lsi  LucidaSans-Italic
hlsr8a  lsr  LucidaSans
hlsb8a  lsb  LucidaSans-Bold
hlsbi8a lsbi  LucidaSans-BoldItalic
hlcrf8a lbl  LucidaBlackletter
hlcriw8a lbh  LucidaHandwriting-Italic
hlcrie8a lbc  LucidaCalligraphy-Italic
hlcrn8a lbkr  LucidaCasual
hlcrin8a lbki  LucidaCasual-Italic
hlsrt8a lstr  LucidaSans-Typewriter
hlsrot8a lsto  LucidaSans-TypewriterOblique
hlsbot8a lstbo LucidaSans-TypewriterBoldOblique
hlsbt8a lstb  LucidaSans-TypewriterBold
hlcrt8a lbtr  LucidaTypewriter
hlcbt8a lbtb  LucidaTypewriterBold
hlcbot8a lbto  LucidaTypewriterOblique
hlcrot8a lbtbo LucidaTypewriterBoldOblique
hlcra  lbma LucidaNewMath-Arrows
hlcda  lbmad LucidaNewMath-Arrows-Demi
hlcrv  lbme LucidaNewMath-Extension
hlcry  lbms LucidaNewMath-Symbol
hlcdy  lbmsd LucidaNewMath-Symbol-Demi
hlcrim  lbmi LucidaNewMath-Italic
hlcrima lbmo LucidaNewMath-AltItalic
hlcdim  lbmdi LucidaNewMath-DemiItalic
hlcdima lbmdo LucidaNewMath-AltDemiItalic
hlcrm  lbmr LucidaNewMath-Roman
hlcdm  lbmd LucidaNewMath-Demibold

3

2 Font description files

2.1 Maths fonts font description files

1 ⟨∗lucidascale⟩

 2 \@ifundefined{DeclareLucidaFontShape}{%
 3 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
 4 \DeclareFontShape{#1}{#2}{#3}{#4}{<->#5}{#6}}}{}

5 ⟨/lucidascale⟩ 6 ⟨∗OMShlh⟩

 7 \ProvidesFile{OMShlh.fd} [1995/02/19]
 8 \DeclareFontFamily{OMS}{hlh}{\skewchar\font'60}
 9 \DeclareFontShape{OMS}{hlh}{m}{n}    {<-> ssub * hlcy/m/n}{}
10 \DeclareFontShape{OMS}{hlh}{m}{it}   {<-> ssub * hlcy/m/n}{}
11 \DeclareFontShape{OMS}{hlh}{m}{sl}   {<-> ssub * hlcy/m/n}{}
12 \DeclareFontShape{OMS}{hlh}{m}{sc}   {<-> ssub * hlcy/m/n}{}
13 \DeclareFontShape{OMS}{hlh}{bx}{n}   {<-> ssub * hlcy/b/n}{}
14 \DeclareFontShape{OMS}{hlh}{bx}{it}  {<-> ssub * hlcy/b/n}{}
15 \DeclareFontShape{OMS}{hlh}{bx}{sl}  {<-> ssub * hlcy/b/n}{}
16 \DeclareFontShape{OMS}{hlh}{bx}{sc}  {<-> ssub * hlcy/b/n}{}

17 ⟨/OMShlh⟩ 18 ⟨∗OMLhlh⟩

19 \ProvidesFile{OMLhlh.fd} [1995/02/19]
20 \DeclareFontShape{OML}{hlh}{m}{n}    {<-> ssub * hlcm/m/n}{}
21 \DeclareFontShape{OML}{hlh}{m}{it}   {<-> ssub * hlcm/m/n}{}
22 \DeclareFontShape{OML}{hlh}{m}{sl}   {<-> ssub * hlcm/m/n}{}
23 \DeclareFontShape{OML}{hlh}{m}{sc}   {<-> ssub * hlcm/m/n}{}
24 \DeclareFontShape{OML}{hlh}{bx}{n}   {<-> ssub * hlcm/b/n}{}
25 \DeclareFontShape{OML}{hlh}{bx}{it}  {<-> ssub * hlcm/b/n}{}
26 \DeclareFontShape{OML}{hlh}{bx}{sl}  {<-> ssub * hlcm/b/n}{}
27 \DeclareFontShape{OML}{hlh}{bx}{sc}  {<-> ssub * hlcm/b/n}{}

28 ⟨/OMLhlh⟩ 29 ⟨∗OMLhlcm⟩

30 % Math Italics
31 \DeclareFontFamily{OML}{hlcm}{\skewchar\font=127}
32 \DeclareLucidaFontShape{OML}{hlcm}{m}{n}{hlcrm}{}
33 \DeclareLucidaFontShape{OML}{hlcm}{b}{n}{hlcdm}{}
34 \DeclareLucidaFontShape{OML}{hlcm}{m}{it}{hlcrima}{}
35 \DeclareLucidaFontShape{OML}{hlcm}{b}{it}{hlcdima}{}

36 ⟨/OMLhlcm⟩ 37 ⟨∗OMShlcy⟩

38 % LucidaNewMath-Symbols
39 \DeclareFontFamily{OMS}{hlcy}{\skewchar\font=48}
40 \DeclareLucidaFontShape{OMS}{hlcy}{m}{n}{hlcry}{}
41 \DeclareLucidaFontShape{OMS}{hlcy}{b}{n}{hlcdy}{}

42 ⟨/OMShlcy⟩ 43 ⟨∗OMXhlcv⟩

44 % LucidaNewMath-Extension
45 \DeclareFontFamily{OMX}{hlcv}{}
46 \DeclareLucidaFontShape{OMX}{hlcv}{m}{n}{hlcrv}{}

47 ⟨/OMXhlcv⟩ 48 ⟨∗LMRhlcm⟩

49 % LucidaNewMath-Arrows
50 \DeclareFontFamily{LMR}{hlcm}{}
51 \DeclareLucidaFontShape{LMR}{hlcm}{m}{n}{hlcra}{}
52 \DeclareLucidaFontShape{LMR}{hlcm}{b}{n}{hlcda}{}

53 ⟨/LMRhlcm⟩ 54 ⟨∗OMLplcm⟩

55 % Math Italics
56 \DeclareFontFamily{OML}{plcm}{\skewchar\font=127}
57 \DeclareFontShape{OML}{plcm}{m}{it}{<->plcm}{}

58 ⟨/OMLplcm⟩ 59 ⟨∗OMSplcy⟩

60 % Math-Symbols
61 \DeclareFontFamily{OMS}{plcy}{\skewchar\font=48}
62 \DeclareFontShape{OMS}{plcy}{m}{n}{<->plcy}{}

63 ⟨/OMSplcy⟩ 64 ⟨∗OMXplcv⟩

65 % Math-Extension
66 \DeclareFontFamily{OMX}{plcv}{}
67 \DeclareFontShape{OMX}{plcv}{m}{n}{<->plcv}{}

68 ⟨/OMXplcv⟩

2.2 (Y&Y names) Lucida Bright font description files

69 ⟨∗OT1lb⟩

70 \typeout{File\space OT1lb.fd\space loading \space (Y\&Y names)\space Lucida Bright}%
71 \DeclareFontFamily{OT1}{lb}{}%
72 \DeclareLucidaFontShape{OT1}{lb}{m}{n}{lbr}{}%
73 \DeclareLucidaFontShape{OT1}{lb}{m}{it}{lbi}{}%
74 \DeclareLucidaFontShape{OT1}{lb}{m}{sl}{lbsl}{}%
75 \DeclareLucidaFontShape{OT1}{lb}{m}{sc}{lbrsc}{}%
76 \DeclareLucidaFontShape{OT1}{lb}{b}{n}{lbd}{}%
77 \DeclareLucidaFontShape{OT1}{lb}{b}{it}{lbdi}{}%
78 \DeclareLucidaFontShape{OT1}{lb}{b}{sc}{lbdsc}{}%
79 \DeclareFontShape{OT1}{lb}{b}{sl}{<->ssub * lb/b/it}{}%
80 \DeclareFontShape{OT1}{lb}{bx}{n}{<->ssub * lb/b/n}{}%
81 \DeclareFontShape{OT1}{lb}{bx}{it}{<->ssub * lb/b/it}{}%
82 \DeclareFontShape{OT1}{lb}{bx}{sl}{<->ssub * lb/b/sl}{}%
83 \DeclareFontShape{OT1}{lb}{bx}{sc}{<->ssub * lb/b/sc}{}%

84 ⟨/OT1lb⟩

2.3 (Y&Y names) Lucida Sans font description files

85 ⟨∗OT1lbs⟩

86 \typeout{File\space OT1lbs.fd\space loading \space (Y\&Y names)\space Lucida Sans}%
87 \DeclareFontFamily{OT1}{lbs}{}%
88 \DeclareLucidaFontShape{OT1}{lbs}{m}{n}{lsr}{}%
89 \DeclareLucidaFontShape{OT1}{lbs}{m}{it}{lsi}{}%

 90 \DeclareLucidaFontShape{OT1}{lbs}{b}{n}{lsd}{}%
 91 \DeclareLucidaFontShape{OT1}{lbs}{b}{it}{lsdi}{}%
 92 \DeclareFontShape{OT1}{lbs}{m}{sl}{<->ssub * lbs/m/it}{}%
 93 \DeclareFontShape{OT1}{lbs}{m}{sc}{<->ssub * lbs/m/n}{}%
 94 \DeclareFontShape{OT1}{lbs}{b}{sc}{<->ssub * lbs/m/sc}{}%
 95 \DeclareFontShape{OT1}{lbs}{bx}{sc}{<->ssub * lbs/b/sc}{}%
 96 \DeclareFontShape{OT1}{lbs}{b}{sl}{<->ssub * lbs/b/it}{}%
 97 \DeclareFontShape{OT1}{lbs}{bx}{n}{<->ssub * lbs/b/n}{}%
 98 \DeclareFontShape{OT1}{lbs}{bx}{it}{<->ssub * lbs/b/it}{}%
 99 \DeclareFontShape{OT1}{lbs}{bx}{sl}{<->ssub * lbs/m/sl}{}%

100 ⟨/OT1lbs⟩

2.4 (Y&Y names) Lucida Bright Typewriter font description files

101 ⟨∗OT1lbt⟩

102 \typeout{File\space OT1lbt.fd\space loading \space (Y\&Y names)\space Lucida Bright Typewriter}%
103 \DeclareFontFamily{OT1}{lbt}{\hyphenchar \font\m@ne}%
104 \DeclareLucidaFontShape{OT1}{lbt}{m}{n}{lstr}{}%
105 \DeclareFontShape{OT1}{lbt}{m}{sc}{<->ssub * lbt/m/n}{}%
106 \DeclareLucidaFontShape{OT1}{lbt}{m}{it}{lsto}{}%
107 \DeclareFontShape{OT1}{lbt}{m}{sl}{<->ssub * lbt/m/it}{}%
108 \DeclareLucidaFontShape{OT1}{lbt}{b}{n}{lstb}{}%
109 \DeclareLucidaFontShape{OT1}{lbt}{b}{it}{lstbo}{}%
110 \DeclareFontShape{OT1}{lbt}{b}{sc}{<->ssub * lbt/m/sc}{}%
111 \DeclareFontShape{OT1}{lbt}{b}{sl}{<->ssub * lbt/b/it}{}%
112 \DeclareFontShape{OT1}{lbt}{m}{sl}{<->ssub * lbt/m/it}{}%
113 \DeclareFontShape{OT1}{lbt}{bx}{n}{<->ssub * lbt/b/n}{}%
114 \DeclareFontShape{OT1}{lbt}{bx}{it}{<->ssub * lbt/b/it}{}%
115 \DeclareFontShape{OT1}{lbt}{bx}{sc}{<->ssub * lbt/b/sc}{}%
116 \DeclareFontShape{OT1}{lbt}{bx}{sl}{<->ssub * lbt/m/sl}{}%

117 ⟨/OT1lbt⟩

2.5 (Y&Y names) Lucida Fax font description files

118 ⟨∗OT1lbf⟩

119 \typeout{File\space OT1lbf.fd\space loading \space (Y\&Y names)\space Lucida Fax}%
120 \DeclareFontFamily{OT1}{lbf}{}%
121 \DeclareLucidaFontShape{OT1}{lbf}{m}{n}{lfr}{}%
122 \DeclareLucidaFontShape{OT1}{lbf}{b}{n}{lfd}{}%
123 \DeclareLucidaFontShape{OT1}{lbf}{m}{it}{lfi}{}%
124 \DeclareLucidaFontShape{OT1}{lbf}{b}{it}{lfdi}{}%

125 ⟨/OT1lbf⟩ 126 ⟨∗OMLlbm⟩

127 % \subsection{Lucida Math Italics  font description files}
128 \DeclareFontFamily{OML}{lbm}{\skewchar\font=127}
129 \DeclareLucidaFontShape{OML}{lbm}{m}{n}{lbmr}{}
130 \DeclareLucidaFontShape{OML}{lbm}{b}{n}{lbmd}{}
131 \DeclareLucidaFontShape{OML}{lbm}{m}{it}{lbmo}{}

132 \DeclareLucidaFontShape{OML}{lbm}{b}{it}{lbmdo}{}

133 ⟨/OMLlbm⟩ 134 ⟨∗OMLlb⟩

135 \ProvidesFile{OMLlb.fd} [1995/02/19]
136 \DeclareFontShape{OML}{lb}{m}{n}    {<-> ssub * lbm/m/n}{}
137 \DeclareFontShape{OML}{lb}{m}{it}   {<-> ssub * lbm/m/n}{}
138 \DeclareFontShape{OML}{lb}{m}{sl}   {<-> ssub * lbm/m/n}{}
139 \DeclareFontShape{OML}{lb}{m}{sc}   {<-> ssub * lbm/m/n}{}
140 \DeclareFontShape{OML}{lb}{bx}{n}   {<-> ssub * lbm/b/n}{}
141 \DeclareFontShape{OML}{lb}{bx}{it}  {<-> ssub * lbm/b/n}{}
142 \DeclareFontShape{OML}{lb}{bx}{sl}  {<-> ssub * lbm/b/n}{}
143 \DeclareFontShape{OML}{lb}{bx}{sc}  {<-> ssub * lbm/b/n}{}

144 ⟨/OMLlb⟩ 145 ⟨∗OMSlb⟩

146 \ProvidesFile{OMSlb.fd} [1995/02/19]
147 \DeclareFontFamily{OMS}{lb}{\skewchar\font'60}
148 \DeclareFontShape{OMS}{lb}{m}{n}    {<-> ssub * lby/m/n}{}
149 \DeclareFontShape{OMS}{lb}{m}{it}   {<-> ssub * lby/m/n}{}
150 \DeclareFontShape{OMS}{lb}{m}{sl}   {<-> ssub * lby/m/n}{}
151 \DeclareFontShape{OMS}{lb}{m}{sc}   {<-> ssub * lby/m/n}{}
152 \DeclareFontShape{OMS}{lb}{bx}{n}   {<-> ssub * lby/b/n}{}
153 \DeclareFontShape{OMS}{lb}{bx}{it}  {<-> ssub * lby/b/n}{}
154 \DeclareFontShape{OMS}{lb}{bx}{sl}  {<-> ssub * lby/b/n}{}
155 \DeclareFontShape{OMS}{lb}{bx}{sc}  {<-> ssub * lby/b/n}{}

156 ⟨/OMSlb⟩ 157 ⟨∗OMSlby⟩

158 % \subsection{LucidaNewMath-Symbols  font description files}
159 \DeclareFontFamily{OMS}{lby}{\skewchar\font=48}
160 \DeclareLucidaFontShape{OMS}{lby}{m}{n}{lbms}{}
161 \DeclareLucidaFontShape{OMS}{lby}{b}{n}{lbmsd}{}

162 ⟨/OMSlby⟩ 163 ⟨∗OMXlbv⟩

164 % \subsection{LucidaNewMath-Extension  font description files}
165 \DeclareFontFamily{OMX}{lbv}{}
166 \DeclareLucidaFontShape{OMX}{lbv}{m}{n}{lbme}{}

167 ⟨/OMXlbv⟩ 168 ⟨∗LMRlbm⟩

169 % \subsection{LucidaNewMath-Arrows  font description files}
170 \DeclareFontFamily{LMR}{lbm}{}
171 \DeclareLucidaFontShape{LMR}{lbm}{m}{n}{lbma}{}
172 \DeclareLucidaFontShape{LMR}{lbm}{b}{n}{lbmad}{}

173 ⟨/LMRlbm⟩

3 Packages

174 ⟨∗lucid⟩

175 \ProvidesPackage{lucid}[\filedate\space\fileversion\space
176 Lucida PSNFSS2e package]
177 \renewcommand{\sfdefault}{plcs}

178 \renewcommand{\rmdefault}{plc}
179 \renewcommand{\ttdefault}{pcr}

180 ⟨/lucid⟩ 181 ⟨∗lucmath⟩

182 \ProvidesPackage{lucmath}[\filedate\space\fileversion\space
183 Lucida Math LaTeX2e package]
184 \DeclareSymbolFont{operators}{\encodingdefault}{\rmdefault}{m}{n}
185 \DeclareSymbolFont{letters}{OML}{plcm}{m}{it}
186 \DeclareSymbolFont{symbols}{OMS}{plcy}{m}{n}
187 \DeclareSymbolFont{largesymbols}{OMX}{plcv}{m}{n}
188 \DeclareSymbolFont{italics}{\encodingdefault}{\rmdefault}{m}{it}
189 \SetSymbolFont{letters}{bold}{OML}{plcm}{m}{it}%
190 \SetSymbolFont{operators}{bold}{\encodingdefault}{\rmdefault}{b}{n}%
191 \SetSymbolFont{operators}{normal}{\encodingdefault}{\rmdefault}{m}{n}%
192 % reset the text in math macros

193 %

194 \DeclareMathAlphabet      {\mathbf}{\encodingdefault}{\rmdefault}{b}{n}%
195 \DeclareMathAlphabet      {\mathrm}{\encodingdefault}{\rmdefault}{m}{n}%
196 \DeclareMathAlphabet      {\mathsf}{\encodingdefault}{\sfdefault}{m}{n}%
197 \DeclareMathAlphabet      {\mathit}{\encodingdefault}{\rmdefault}{m}{it}%
198 \DeclareMathAlphabet      {\mathtt}{\encodingdefault}{\ttdefault}{m}{n}%
199 \SetMathAlphabet{\mathbf}{bold}{\encodingdefault}{\rmdefault}{b}{n}%
200 \SetMathAlphabet{\mathsf}{bold}{\encodingdefault}{\sfdefault}{b}{n}%
201 \SetMathAlphabet{\mathrm}{bold}{\encodingdefault}{\rmdefault}{b}{n}%
202 \SetMathAlphabet{\mathit}{bold}{\encodingdefault}{\rmdefault}{b}{it}%
203 \SetMathAlphabet{\mathtt}{bold}{\encodingdefault}{\ttdefault}{b}{n}%

204 ⟨/lucmath⟩ 205 ⟨∗lucmtime⟩

206 \ProvidesPackage{lucmtime}[\filedate\space\fileversion\space
207 Monotype Times fonts with Lucida New Math]
208 \def\rmdefault{mntx}
209 \def\sfdefault{cmss}
210 \def\ttdefault{cmtt}
211 \def\Mathdefault{mntluc}
212 \DeclareSymbolFont{letters}{OML}{mntluc}{m}{it}
213 \DeclareSymbolFont{operators}{OT1}{mntx}{m}{n}
214 \SetSymbolFont{letters}{normal}{OML}{mntluc}{m}{it}
215 \SetSymbolFont{letters}{bold}{OML}{mntluc}{b}{it}
216 \SetSymbolFont{operators}{bold}{OT1}{mntx}{b}{n}%
217 \SetSymbolFont{operators}{normal}{OT1}{mntx}{m}{n}%

218 ⟨/lucmtime⟩ 219 ⟨∗lucmin⟩

220 \ProvidesPackage{lucmin}[\filedate\space\fileversion\space
221 Minion fonts with Lucida New Math]
222 \def\rmdefault{zmn}
223 \def\sfdefault{zmy}
224 \def\ttdefault{hlct}
225 \renewcommand{\bfdefault}{b}
226 \def\Mathdefault{zmnluc}
227 \DeclareSymbolFont{letters}{OML}{zmnluc}{m}{it}

228 \DeclareSymbolFont{operators}{OT1}{zmn}{m}{n}
229 \SetSymbolFont{letters}{normal}{OML}{zmnluc}{m}{it}
230 \SetSymbolFont{letters}{bold}{OML}{zmnluc}{b}{it}
231 \SetSymbolFont{operators}{bold}{OT1}{zmn}{b}{n}%
232 \SetSymbolFont{operators}{normal}{OT1}{zmn}{m}{n}%

233 ⟨/lucmin⟩ 234 ⟨∗lucidabright⟩

235 \ProvidesPackage{lucbr}[\filedate\space\fileversion\space
236 Lucida Bright PSNFSS2e package]

237 ⟨/lucidabright⟩ 238 ⟨∗lucidabright | lucbmath⟩

239 \newif\iflucida@expert
240 \DeclareOption{kb}{% this is for Berry-type font names
241 \def\Lucida@names{0}}
242 \DeclareOption{yy}{% this is for Y\&Y font names
243 \def\Lucida@names{1}}
244 \DeclareOption{expert}{% we have expert set
245 \lucida@experttrue}
246 \DeclareOption{noexpert}{% we do not have expert set
247 \lucida@expertfalse}

Set up the variant text and math sizes which Y&Y suggest for Lucida. The figures for these two options actually come from Frank Mittelbach (oh great one).

The default is to scale, but two options allow you to revert to normal behaviour, or get even smaller.

248 \DeclareOption{nolucidascale}{%
249 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
250  \DeclareFontShape{#1}{#2}{#3}{#4}{<->#5}{#6}%

251 }% 252 }

253 \DeclareOption{lucidascale}{%
254 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
255 \DeclareFontShape{#1}{#2}{#3}{#4}{%
256 <-5.5>s*[1.04]#5%
257 <5.5-6.5>s*[1.02]#5%
258 <6.5-7.5>s*[.99]#5%
259 <7.5-8.5>s*[.97]#5%
260 <8.5-9.5>s*[.96]#5%
261 <9.5-10.5>s*[.95]#5%
262 <10.5-11.5>s*[.94]#5%
263 <11.5-13>s*[.93]#5%
264 <13-15.5>s*[.92]#5%
265 <15.5-18.5>s*[.91]#5%
266 <18.5-22.5>s*[.9]#5%
267 <22.5->s*[.89]#5%
268 }{#6}}

269 }

270 \DeclareOption{lucidasmallscale}{%
271 \def\DeclareLucidaFontShape#1#2#3#4#5#6{%
272 \DeclareFontShape{#1}{#2}{#3}{#4}{%

273 <-5.5>s*[.98]#5%
274 <5.5-6.5>s*[.96]#5%
275 <6.5-7.5>s*[.94]#5%
276 <7.5-8.5>s*[.92]#5%
277 <8.5-9.5>s*[.91]#5%
278 <9.5-10.5>s*[.9]#5%
279 <10.5-11.5>s*[.89]#5%
280 <11.5-13>s*[.88]#5%
281 <13-15.5>s*[.87]#5%
282 <15.5-18.5>s*[.86]#5%
283 <18.5-22.5>s*[.85]#5%
284 <22.5->s*[.84]#5%
285 }{#6}}

286 }

The same effect as some of the oldlfont package; in particular, its allows \rm, \bf etc in maths.

287 \DeclareOption{oldlfont}{%
288 \AtBeginDocument{%
289 \let\math@bgroup\@empty
290 \let\math@egroup\@empty
291 \let \@@math@bgroup \math@bgroup
292 \let \@@math@egroup \math@egroup
293 \let\mit\undefined
294 \let\cal\undefined
295 \DeclareSymbolFontAlphabet{\mit}{letters}%
296 \DeclareSymbolFontAlphabet{\cal}{symbols}%
297 \DeclareRobustCommand\rm{\normalfont\rmfamily\mathgroup\symoperators}%
298 \DeclareRobustCommand\bf{\normalfont\bfseries\mathgroup\symbold}%
299 \DeclareRobustCommand\it{\normalfont\itshape\mathgroup\symitalic}%
300 \DeclareRobustCommand\tt{\normalfont\ttfamily\mathgroup\symtypewriter}%
301 \DeclareRobustCommand\em{%
302 \@nomath\em
303 \ifdim \fontdimen\@ne\font>\z@\rm\else\it\fi}%
304 \def\@setfontsize##1##2##3{\@nomath##1%
305  \ifx\protect\@typeset@protect
306  \let\@currsize##1%
307  \fi
308  \fontsize{##2}{##3}\normalfont}%
309 \let\not@math@alphabet\@gobbletwo

310 } 311 }

312 \ExecuteOptions{kb,expert,lucidascale}
313 \ProcessOptions

314 ⟨/lucidabright | lucbmath⟩ 315 ⟨lucidabright | lucbmath⟩\ifnum\Lucida@names=0 316 ⟨∗lucidabright⟩

317 \renewcommand{\sfdefault}{hls}
318 \renewcommand{\rmdefault}{hlh}
319 \renewcommand{\ttdefault}{hlst}

320 ⟨/lucidabright⟩ 321 ⟨∗lucbmath⟩

New encoding scheme for Math Arrows font

322 \DeclareFontEncoding{LMR}{}{}
323 \DeclareFontSubstitution{LMR}{hlcm}{m}{n}

324 ⟨!luctim⟩ \DeclareSymbolFont{letters}{OML}{hlcm}{m}{it}

325 \iflucida@expert
326 \DeclareSymbolFont{mathupright}{OML}{hlcm}{m}{n}
327 \fi
328 \DeclareSymbolFont{symbols}{OMS}{hlcy}{m}{n}
329 \DeclareSymbolFont{largesymbols}{OMX}{hlcv}{m}{n}

The new Expert set for bold math

330 \iflucida@expert

331 ⟨!luctim⟩ \SetSymbolFont{letters}{bold}{OML}{hlcm}{b}{it}

332 \SetSymbolFont{mathupright}{bold}{OML}{hlcm}{b}{n}
333 \SetSymbolFont{symbols}{bold}{OMS}{hlcy}{b}{n}
334 \fi

Better get the order of this right, or maths come out as all arrows…

335 \DeclareSymbolFont{italics}{\encodingdefault}{\rmdefault}{m}{it}
336 \DeclareSymbolFont{arrows}{LMR}{hlcm}{m}{n}
337 \iflucida@expert
338 \DeclareSymbolFont{boldarrows}{LMR}{hlcm}{b}{n}
339 \fi

340 ⟨/lucbmath⟩ 341 ⟨lucidabright | lucbmath⟩\else 342 ⟨∗lucidabright⟩

This is for Y&Y font names; the supplied tfms are in OT1 encoding

343 \renewcommand{\sfdefault}{lbs}
344 \renewcommand{\rmdefault}{lb}
345 \renewcommand{\ttdefault}{lbt}
346 \def\encodingdefault{OT1}\fontencoding{OT1}\selectfont

347 ⟨/lucidabright⟩ 348 ⟨∗lucbmath⟩

New encoding scheme for Math Arrows font

349 \DeclareFontEncoding{LMR}{}{}
350 \DeclareFontSubstitution{LMR}{lbm}{m}{n}

351 ⟨!luctim⟩ \DeclareSymbolFont{letters}{OML}{lbm}{m}{it}

352 \iflucida@expert
353 \DeclareSymbolFont{mathupright}{OML}{lbm}{m}{n}
354 \fi
355 \DeclareSymbolFont{symbols}{OMS}{lby}{m}{n}
356 \DeclareSymbolFont{largesymbols}{OMX}{lbv}{m}{n}
357 \iflucida@expert

358 ⟨!luctim⟩ \SetSymbolFont{letters}{bold}{OML}{lbm}{b}{it}

359 \SetSymbolFont{mathupright}{bold}{OML}{lbm}{b}{n}
360 \SetSymbolFont{symbols}{bold}{OMS}{lby}{b}{n}
361 \fi

362 \DeclareSymbolFont{italics}{\encodingdefault}{\rmdefault}{m}{it}
363 \DeclareSymbolFont{arrows}{LMR}{lbm}{m}{n}
364 \iflucida@expert
365 \DeclareSymbolFont{boldarrows}{LMR}{lbm}{b}{n}
366 \fi

367 ⟨/lucbmath⟩ 368 ⟨lucidabright | lucbmath⟩\fi 369 ⟨∗lucbmath⟩ 370 ⟨!luctim⟩\DeclareSymbolFont{operators}{OT1}{\rmdefault}{m}{n} 371 ⟨!luctim⟩\SetSymbolFont{operators}{bold}{OT1}{\rmdefault}{b}{n}% 372 ⟨!luctim⟩\SetSymbolFont{operators}{normal}{OT1}{\rmdefault}{m}{n}%

Explicitly redeclare all the alphabets just in case, but differentiate between pure Lucida, and the Times mixture, since those have genuine OT1 mimics.

373 ⟨∗!luctim⟩ 374 \DeclareMathAlphabet {\mathbf}{\encodingdefault}{\rmdefault}{b}{n}% 375 \DeclareMathAlphabet {\mathrm}{\encodingdefault}{\rmdefault}{m}{n}% 376 \DeclareMathAlphabet {\mathsf}{\encodingdefault}{\sfdefault}{m}{n}% 377 \DeclareMathAlphabet {\mathit}{\encodingdefault}{\rmdefault}{m}{it}% 378 \DeclareMathAlphabet {\mathtt}{\encodingdefault}{\ttdefault}{m}{n}% 379 \SetMathAlphabet{\mathbf}{bold}{\encodingdefault}{\rmdefault}{b}{n}% 380 \SetMathAlphabet{\mathsf}{bold}{\encodingdefault}{\sfdefault}{b}{n}% 381 \SetMathAlphabet{\mathrm}{bold}{\encodingdefault}{\rmdefault}{b}{n}% 382 \SetMathAlphabet{\mathit}{bold}{\encodingdefault}{\rmdefault}{b}{it}% 383 \SetMathAlphabet{\mathtt}{bold}{\encodingdefault}{\ttdefault}{b}{n}% 384 ⟨/!luctim⟩ 385 ⟨∗luctim⟩ 386 \DeclareMathAlphabet {\mathbf}{OT1}{\Mathdefault}{b}{n}% 387 \DeclareMathAlphabet {\mathrm}{OT1}{\Mathdefault}{m}{n}% 388 \DeclareMathAlphabet {\mathsf}{OT1}{\sfdefault}{m}{n}% 389 \DeclareMathAlphabet {\mathit}{OT1}{\Mathdefault}{m}{it}% 390 \DeclareMathAlphabet {\mathtt}{OT1}{\ttdefault}{m}{n}% 391 \SetMathAlphabet{\mathbf}{bold}{OT1}{\Mathdefault}{b}{n}% 392 \SetMathAlphabet{\mathsf}{bold}{OT1}{\sfdefault}{b}{n}% 393 \SetMathAlphabet{\mathrm}{bold}{OT1}{\Mathdefault}{b}{n}% 394 \SetMathAlphabet{\mathit}{bold}{OT1}{\Mathdefault}{b}{it}% 395 \SetMathAlphabet{\mathtt}{bold}{OT1}{\ttdefault}{b}{n}% 396 ⟨/luctim⟩

397 \DeclareSymbolFontAlphabet{\mathbb}{arrows}
398 \DeclareSymbolFontAlphabet{\mathscr}{symbols}
399 \iflucida@expert
400 \DeclareSymbolFontAlphabet{\mathup}{mathupright}
401 \fi
402 \ifnum\Lucida@names=0
403 \DeclareMathAccent\vec  {\mathord}{letters}{"7E}
404 \else
405 \DeclareMathAccent\vec  {\mathord}{letters}{"7E}
406 \DeclareMathAccent\dot  {\mathalpha}{operators}{"C7}
407 \fi

This section derives mostly from Berthold Horn’s lcdmacro.tex and amssymblb.tex

\begin{aligned}& c \\ ◯ \end{aligned}

1991, 1992 Y&Y. All Rights Reserved Original from Version 1.2, 1992 June 14; updated ad hoc.

408 \@ifpackageloaded{amsmath}{%

(From M J Downes): it’s possible the factors 1.5, 2, 2.5, 3, 3.5 should be adjusted for Lucida fonts. But that has to be determined by looking at printed tests which I cannot do at the moment. [mjd,24-Jun-1993]

409 \def\biggg{\bBigg@\thr@@}
410 \def\Biggg{\bBigg@{3.5}}
411 }{%
412 \def\big#1{{\hbox{$\left#1\vbox to8.20\p@{}\right.\n@space$}}}
413 \def\Big#1{{\hbox{$\left#1\vbox to10.80\p@{}\right.\n@space$}}}
414 \def\bigg#1{{\hbox{$\left#1\vbox to13.42\p@{}\right.\n@space$}}}
415 \def\Bigg#1{{\hbox{$\left#1\vbox to16.03\p@{}\right.\n@space$}}}
416 \def\biggg#1{{\hbox{$\left#1\vbox to17.72\p@{}\right.\n@space$}}}
417 \def\Biggg#1{{\hbox{$\left#1\vbox to21.25\p@{}\right.\n@space$}}}
418 \def\n@space{\nulldelimiterspace\z@ \m@th}

419 }

Define some extra large sizes — always done using extensible parts

420 \def\bigggl{\mathopen\biggg}
421 \def\bigggr{\mathclose\biggg}
422 \def\Bigggl{\mathopen\Biggg}
423 \def\Bigggr{\mathclose\Biggg}

Following is only really needed if the roman text font is NOT LucidaBright Draw the small sizes of ‘[’ and ‘]’ from math italic instead of roman font

424 \mathcode`\[="4186 \delcode`\[="186302
425 \mathcode`\]="5187 \delcode`\]="187303

Draw the small sizes of ‘(’ and ‘)’ from math italic instead of roman font

426 \mathcode`\(="4184 \delcode`\(="184300
427 \mathcode`\)="5185 \delcode`\)="185301

Draw ‘=’ and ‘+’ from symbol font instead of roman

428 \mathcode`\=="3283
429 \mathcode`\+="2282

Draw small ‘/’ from math italic instead of roman font

430 \mathcode`\/="013D \delcode`\/="13D30E

Make open face brackets accessible, i.e. [[ and ]]

431 \def\ldbrack{\delimiter"4182382 }
432 \def\rdbrack{\delimiter"5183383 }

Provide access to surface integral signs (linked from text to display size)

433 \DeclareMathSymbol{\surfintop}{1}{largesymbols}{"90}
434 \def\surfint{\surfintop\nolimits}

Make medium size integrals available (NOT linked to display size)

435 \DeclareMathSymbol{\midintop}{1}{largesymbols}{"92}

436 \def\midint{\midintop\nolimits}
437 \DeclareMathSymbol{\midointop}{1}{largesymbols}{"93}
438 \def\midoint{\midointop\nolimits}
439 \DeclareMathSymbol{\midsurfintop}{1}{largesymbols}{"94}
440 \def\midsurfint{\midsurfintop\nolimits}

Extensible integral (use with \bigg, \Bigg, \biggg, \Biggg etc)

441 \def\largeint{\delimiter"135A395 }

To close up gaps in special math characters constructed from pieces

442 \def\joinrel{\mathrel{\mkern-4mu}} % \def\joinrel{\mathrel{\mkern-3mu}}

The \mkern-2.5mu undoes the bogus ‘italic correction’ after joiners in LBMA

443 \def\relbar{\mathrel{\smash{\mathchar"3\hexnumber@\symarrows 2D}}\mathrel{\mkern-2.5mu}}
444 \def\Relbar{\mathrel{\mathchar"3\hexnumber@\symarrows 3D}\mathrel{\mkern-2.5mu}}

The \mkern4mu undoes the overhang at the ends of the joiners (and more)

445 \def\longleftarrow{\leftarrow\relbar\mathrel{\mkern4mu}}
446 \def\longrightarrow{\mathrel{\mkern4mu}\relbar\rightarrow}
447 \def\Longleftarrow{\Leftarrow\Relbar\mathrel{\mkern4mu}}
448 \def\Longrightarrow{\mathrel{\mkern4mu}\Relbar\Rightarrow}

Some characters that need construction in CM exist complete in math italic or math symbol font.

449 \let\bowtie\undefined
450 \let\models\undefined
451 \let\doteq\undefined
452 \let\cong\undefined
453 \let\angle\undefined
454 \DeclareMathSymbol{\bowtie}{3}{letters}{"F6}
455 \DeclareMathSymbol{\models}{3}{symbols}{"EE}
456 \DeclareMathSymbol{\doteq}{3}{symbols}{"C9}
457 \DeclareMathSymbol{\cong}{3}{symbols}{"9B}
458 \DeclareMathSymbol{\angle}{0}{symbols}{"8B}

These need undefining so that we can redeclare them.

459 \let\Box\undefined
460 \let\Diamond\undefined
461 \let\leadsto\undefined
462 \let\neq\undefined
463 \let\hookleftarrow\undefined
464 \let\hookrightarrow\undefined
465 \let\mapsto\undefined
466 \let\notin\undefined
467 \let\rightleftharpoons\undefined

Other characters may be found in LucidaNewMath-Arrows (more negated later).

468 \DeclareMathSymbol{\neq}{3}{arrows}{"94}
469 \DeclareMathSymbol{\rightleftharpoons}{3}{arrows}{"7A}
470 \DeclareMathSymbol{\leftrightharpoons}{3}{arrows}{"79}
471 \DeclareMathSymbol{\hookleftarrow}{3}{arrows}{"3C}
472 \DeclareMathSymbol{\hookrightarrow}{3}{arrows}{"3E}

473 \DeclareMathSymbol{\mapsto}{3}{arrows}{"2C}
474 \def\longmapsto{\mapstochar\longrightarrow}

Special L^{A}TEX character definitions (originally from L^{A}TEX symbol font)

475 \let\Join\undefined
476 \let\rhd\undefined
477 \let\lhd\undefined
478 \let\unrhd\undefined
479 \let\unlhd\undefined
480 \DeclareMathSymbol{\Join}{3}{letters}{"F6}
481 \DeclareMathSymbol{\rhd}{3}{letters}{"2E}
482 \DeclareMathSymbol{\lhd}{3}{letters}{"2F}
483 \DeclareMathSymbol{\unlhd}{3}{symbols}{"F4}
484 \DeclareMathSymbol{\unrhd}{3}{symbols}{"F5}
485 \DeclareMathSymbol{\Box}{0}{arrows}{"02}
486 \DeclareMathSymbol{\Diamond}{0}{arrows}{"08}
487 \DeclareMathSymbol{\leadsto}{3}{arrows}{"8E}
488 \DeclareMathSymbol{\leadsfrom}{3}{arrows}{"8D}
489 \def\mathstrut{\vphantom{f}}
490 \@ifpackageloaded{amstex}{}{%
491 \@ifpackageloaded{amsmath}{}{%

AMS math not in use: modify \matrix it to adjust the first and last line vertical spacing slightly; otherwise leave it alone. following changed because fonts (i.e. math italic) not ‘at full scale’

492 \def\matrix#1{\null\,\vcenter{\normalbaselines\m@th
493  \ialign{\hfil$##$\hfil&&\quad\hfil$##$\hfil\crcr
494  \mathstrut\crcr\noalign{\kern-0.9\baselineskip}
495  #1\crcr\mathstrut\crcr\noalign{\kern-0.9\baselineskip}}}\,}

496 } 497 }

In n-th root, don’t want the ‘n’ to come too close to the radical

498 \def\r@@t#1#2{\setbox\z@\hbox{$\m@th#1\sqrt{#2}$}
499 \dimen@\ht\z@ \advance\dimen@-\dp\z@
500 \mkern5mu\raise.6\dimen@\copy\rootbox \mkern-7.5mu \box\z@}

Here are some extra definitions of mathematical symbols and operators:

501 \DeclareMathSymbol{\defineequal}{3}{symbols}{"D6}
502 \DeclareMathSymbol{\notleq}{3}{arrows}{"9C}
503 \DeclareMathSymbol{\notgeq}{3}{arrows}{"9D}
504 \DeclareMathSymbol{\notequiv}{3}{arrows}{"95}
505 \DeclareMathSymbol{\notprec}{3}{arrows}{"E5}
506 \DeclareMathSymbol{\notsucc}{3}{arrows}{"E6}
507 \DeclareMathSymbol{\notapprox}{3}{arrows}{"98}
508 \DeclareMathSymbol{\notpreceq}{3}{arrows}{"E7}
509 \DeclareMathSymbol{\notsucceq}{3}{arrows}{"E8}
510 \DeclareMathSymbol{\notasymp}{3}{arrows}{"F3}
511 \DeclareMathSymbol{\notsubset}{3}{arrows}{"C6}
512 \DeclareMathSymbol{\notsupset}{3}{arrows}{"C7}
513 \DeclareMathSymbol{\notsim}{3}{arrows}{"96}

514 \DeclareMathSymbol{\notsubseteq}{3}{arrows}{"C8}
515 \DeclareMathSymbol{\notsupseteq}{3}{arrows}{"C9}
516 \DeclareMathSymbol{\notsimeq}{3}{arrows}{"97}
517 \DeclareMathSymbol{\notsqsubseteq}{3}{arrows}{"D4}
518 \DeclareMathSymbol{\notsqsupseteq}{3}{arrows}{"D5}
519 \DeclareMathSymbol{\notcong}{3}{arrows}{"99}
520 \DeclareMathSymbol{\notin}{3}{arrows}{"1D}
521 \DeclareMathSymbol{\notni}{3}{arrows}{"1F}
522 \DeclareMathSymbol{\notvdash}{3}{arrows}{"F8}
523 \DeclareMathSymbol{\notmodels}{3}{arrows}{"F9}
524 \DeclareMathSymbol{\notparallel}{3}{arrows}{"F7}
525 \DeclareMathSymbol{\noteq}{3}{arrows}{"94}
526 \DeclareMathSymbol{\notless}{3}{arrows}{"9A}
527 \DeclareMathSymbol{\notgreater}{3}{arrows}{"9B}
528 \DeclareMathSymbol{\notmid}{3}{arrows}{"F6}
529 \let\Bbb\mathbb

Normal L^{A}TEX draws upper case (upright) Greek from cmr10 — when using the Cork encoding, that isn’t there.

530 \iflucida@expert

If we have the LucidaBright Expert set, we’ll draw them from the upright math font. That way we can get bold math to work on upright upper case Greek.

Why doesn’t this work?

  \documentclass{article}
  \usepackage{lucbry}
  $\mathbf{\Sigma}$
  \end{document}

The answer lies in the meaning of \mathbf; as fntguide.tex says, it is for alphabetic switching. The straight lucida style says

   \DeclareMathSymbol{\Sigma}{\mathalpha}{largesymbols}{'326}

and the \mathalpha signifies that the \Sigma can change with the alphabet; so this in fact looks for \char'326 in the “mathbf” alphabet when we ask for that. That is defined with

  \SetMathAlphabet{\mathbf}{bold}{\encodingdefault}{\rmdefault}{b}{n}

ie normal text Lucida bold. It all works in CMR because the text fonts have Greek, which is why the symbols are defined as ; in addition, the alphabets like \mathbf explicitly ask for OT1:

 \DeclareMathAlphabet      {\mathbf}{OT1}{cmr}{bx}{n}

so it works in T1 encoding too.

When we get the symbols from other fonts in Lucida, we should no longer classify the fonts as \mathalpha, since the mechanism doesn’t function. So we use \mathord instead, and you only get bold Greek if you change \mathversion. At least it’s consistent.

If, however, we are using the Times mixture, we can keep \mathalpha, as we have the right font layouts around.

531 ⟨∗!luctim⟩ 532 \DeclareMathSymbol{\Gamma}{\mathord}{mathupright}{0} 533 \DeclareMathSymbol{\Delta}{\mathord}{mathupright}{1} 534 \DeclareMathSymbol{\Theta}{\mathord}{mathupright}{2} 535 \DeclareMathSymbol{\Lambda}{\mathord}{mathupright}{3} 536 \DeclareMathSymbol{\Xi}{\mathord}{mathupright}{4} 537 \DeclareMathSymbol{\Pi}{\mathord}{mathupright}{5} 538 \DeclareMathSymbol{\Sigma}{\mathord}{mathupright}{6} 539 \DeclareMathSymbol{\Upsilon}{\mathord}{mathupright}{7} 540 \DeclareMathSymbol{\Phi}{\mathord}{mathupright}{8} 541 \DeclareMathSymbol{\Psi}{\mathord}{mathupright}{9} 542 \DeclareMathSymbol{\Omega}{\mathord}{mathupright}{10} 543 \DeclareMathSymbol{\upalpha}{\mathord}{mathupright}{11} 544 \DeclareMathSymbol{\upbeta}{\mathord}{mathupright}{12} 545 \DeclareMathSymbol{\upgamma}{\mathord}{mathupright}{13} 546 \DeclareMathSymbol{\updelta}{\mathord}{mathupright}{14} 547 \DeclareMathSymbol{\upepsilon}{\mathord}{mathupright}{15} 548 \DeclareMathSymbol{\upzeta}{\mathord}{mathupright}{16} 549 \DeclareMathSymbol{\upeta}{\mathord}{mathupright}{17} 550 \DeclareMathSymbol{\uptheta}{\mathord}{mathupright}{18} 551 \DeclareMathSymbol{\upiota}{\mathord}{mathupright}{19} 552 \DeclareMathSymbol{\upkappa}{\mathord}{mathupright}{20} 553 \DeclareMathSymbol{\uplambda}{\mathord}{mathupright}{21} 554 \DeclareMathSymbol{\upmu}{\mathord}{mathupright}{22} 555 \DeclareMathSymbol{\upnu}{\mathord}{mathupright}{23} 556 \DeclareMathSymbol{\upxi}{\mathord}{mathupright}{24} 557 \DeclareMathSymbol{\uppi}{\mathord}{mathupright}{25} 558 \DeclareMathSymbol{\uprho}{\mathord}{mathupright}{26} 559 \DeclareMathSymbol{\upsigma}{\mathord}{mathupright}{27} 560 \DeclareMathSymbol{\uptau}{\mathord}{mathupright}{28} 561 \DeclareMathSymbol{\upupsilon}{\mathord}{mathupright}{29} 562 \DeclareMathSymbol{\upphi}{\mathord}{mathupright}{30} 563 \DeclareMathSymbol{\upchi}{\mathord}{mathupright}{31} 564 \DeclareMathSymbol{\uppsi}{\mathord}{mathupright}{32} 565 \DeclareMathSymbol{\upomega}{\mathord}{mathupright}{33} 566 \DeclareMathSymbol{\upvarepsilon}{\mathord}{mathupright}{34} 567 \else

It’s in the extension font (largesymbols)

568 \DeclareMathSymbol{\Gamma}{\mathord}{largesymbols}{’320} 569 \DeclareMathSymbol{\Delta}{\mathord}{largesymbols}{’321} 570 \DeclareMathSymbol{\Theta}{\mathord}{largesymbols}{’322} 571 \DeclareMathSymbol{\Lambda}{\mathord}{largesymbols}{’323} 572 \DeclareMathSymbol{\Xi}{\mathord}{largesymbols}{’324} 573 \DeclareMathSymbol{\Pi}{\mathord}{largesymbols}{’325} 574 \DeclareMathSymbol{\Sigma}{\mathord}{largesymbols}{’326} 575 \DeclareMathSymbol{\Upsilon}{\mathord}{largesymbols}{’327} 576 \DeclareMathSymbol{\Phi}{\mathord}{largesymbols}{’330} 577 \DeclareMathSymbol{\Psi}{\mathord}{largesymbols}{’331} 578 \DeclareMathSymbol{\Omega}{\mathord}{largesymbols}{’332} 579 \fi 580 ⟨/!luctim⟩ 581 ⟨∗luctim⟩ 582 \DeclareMathSymbol{\Gamma}{\mathalpha}{mathupright}{0} 583 \DeclareMathSymbol{\Delta}{\mathalpha}{mathupright}{1} 584 \DeclareMathSymbol{\Theta}{\mathalpha}{mathupright}{2} 585 \DeclareMathSymbol{\Lambda}{\mathalpha}{mathupright}{3} 586 \DeclareMathSymbol{\Xi}{\mathalpha}{mathupright}{4} 587 \DeclareMathSymbol{\Pi}{\mathalpha}{mathupright}{5} 588 \DeclareMathSymbol{\Sigma}{\mathalpha}{mathupright}{6} 589 \DeclareMathSymbol{\Upsilon}{\mathalpha}{mathupright}{7} 590 \DeclareMathSymbol{\Phi}{\mathalpha}{mathupright}{8} 591 \DeclareMathSymbol{\Psi}{\mathalpha}{mathupright}{9} 592 \DeclareMathSymbol{\Omega}{\mathalpha}{mathupright}{10} 593 \DeclareMathSymbol{\upalpha}{\mathalpha}{mathupright}{11} 594 \DeclareMathSymbol{\upbeta}{\mathalpha}{mathupright}{12} 595 \DeclareMathSymbol{\upgamma}{\mathalpha}{mathupright}{13} 596 \DeclareMathSymbol{\updelta}{\mathalpha}{mathupright}{14} 597 \DeclareMathSymbol{\upepsilon}{\mathalpha}{mathupright}{15} 598 \DeclareMathSymbol{\upzeta}{\mathalpha}{mathupright}{16} 599 \DeclareMathSymbol{\upeta}{\mathalpha}{mathupright}{17} 600 \DeclareMathSymbol{\uptheta}{\mathalpha}{mathupright}{18} 601 \DeclareMathSymbol{\upiota}{\mathalpha}{mathupright}{19} 602 \DeclareMathSymbol{\upkappa}{\mathalpha}{mathupright}{20} 603 \DeclareMathSymbol{\uplambda}{\mathalpha}{mathupright}{21} 604 \DeclareMathSymbol{\upmu}{\mathalpha}{mathupright}{22} 605 \DeclareMathSymbol{\upnu}{\mathalpha}{mathupright}{23} 606 \DeclareMathSymbol{\upxi}{\mathalpha}{mathupright}{24} 607 \DeclareMathSymbol{\uppi}{\mathalpha}{mathupright}{25} 608 \DeclareMathSymbol{\uprho}{\mathalpha}{mathupright}{26} 609 \DeclareMathSymbol{\upsigma}{\mathalpha}{mathupright}{27} 610 \DeclareMathSymbol{\uptau}{\mathalpha}{mathupright}{28} 611 \DeclareMathSymbol{\upupsilon}{\mathalpha}{mathupright}{29} 612 \DeclareMathSymbol{\upphi}{\mathalpha}{mathupright}{30} 613 \DeclareMathSymbol{\upchi}{\mathalpha}{mathupright}{31} 614 \DeclareMathSymbol{\uppsi}{\mathalpha}{mathupright}{32} 615 \DeclareMathSymbol{\upomega}{\mathalpha}{mathupright}{33} 616 \DeclareMathSymbol{\upvarepsilon}{\mathalpha}{mathupright}{34} 617 \else

It’s in the extension font (largesymbols)

618 \DeclareMathSymbol{\Gamma}{\mathalpha}{largesymbols}{’320} 619 \DeclareMathSymbol{\Delta}{\mathalpha}{largesymbols}{’321} 620 \DeclareMathSymbol{\Theta}{\mathalpha}{largesymbols}{’322} 621 \DeclareMathSymbol{\Lambda}{\mathalpha}{largesymbols}{’323} 622 \DeclareMathSymbol{\Xi}{\mathalpha}{largesymbols}{’324} 623 \DeclareMathSymbol{\Pi}{\mathalpha}{largesymbols}{’325} 624 \DeclareMathSymbol{\Sigma}{\mathalpha}{largesymbols}{’326} 625 \DeclareMathSymbol{\Upsilon}{\mathalpha}{largesymbols}{’327} 626 \DeclareMathSymbol{\Phi}{\mathalpha}{largesymbols}{’330} 627 \DeclareMathSymbol{\Psi}{\mathalpha}{largesymbols}{’331} 628 \DeclareMathSymbol{\Omega}{\mathalpha}{largesymbols}{’332} 629 \fi 630 ⟨/luctim⟩

631 \DeclareMathSymbol{\varGamma}{\mathalpha}{letters}{"00}
632 \DeclareMathSymbol{\varDelta}{\mathalpha}{letters}{"01}
633 \DeclareMathSymbol{\varTheta}{\mathalpha}{letters}{"02}
634 \DeclareMathSymbol{\varLambda}{\mathalpha}{letters}{"03}
635 \DeclareMathSymbol{\varXi}{\mathalpha}{letters}{"04}
636 \DeclareMathSymbol{\varPi}{\mathalpha}{letters}{"05}
637 \DeclareMathSymbol{\varSigma}{\mathalpha}{letters}{"06}
638 \DeclareMathSymbol{\varUpsilon}{\mathalpha}{letters}{"07}
639 \DeclareMathSymbol{\varPhi}{\mathalpha}{letters}{"08}
640 \DeclareMathSymbol{\varPsi}{\mathalpha}{letters}{"09}
641 \DeclareMathSymbol{\varOmega}{\mathalpha}{letters}{"0A}

Definitions for math symbols and operators (normally found in the AMS symbol fonts) using LucidaNewMath fonts MSAM* equivalents:

642 \DeclareMathSymbol{\boxdot}{2}{symbols}{"ED}
643 \DeclareMathSymbol{\boxplus}{2}{symbols}{"EA}
644 \DeclareMathSymbol{\boxtimes}{2}{symbols}{"EC}
645 \DeclareMathSymbol{\square}{0}{arrows}{"02}
646 \DeclareMathSymbol{\blacksquare}{0}{arrows}{"03}
647 \DeclareMathSymbol{\centerdot}{2}{arrows}{"E1}
648 \DeclareMathSymbol{\lozenge}{0}{arrows}{"08}
649 \DeclareMathSymbol{\blacklozenge}{0}{arrows}{"09}
650 \DeclareMathSymbol{\circlearrowright}{3}{arrows}{"8C}
651 \DeclareMathSymbol{\circlearrowleft}{3}{arrows}{"8B}
652 \DeclareMathSymbol{\rightleftharpoons}{3}{arrows}{"7A}
653 \DeclareMathSymbol{\leftrightharpoons}{3}{arrows}{"79}
654 \DeclareMathSymbol{\boxminus}{2}{symbols}{"EB}
655 \DeclareMathSymbol{\Vdash}{3}{symbols}{"F0}
656 \DeclareMathSymbol{\Vvdash}{3}{letters}{"D3}
657 \DeclareMathSymbol{\vDash}{3}{symbols}{"EE}
658 \DeclareMathSymbol{\twoheadrightarrow}{3}{arrows}{"25}
659 \DeclareMathSymbol{\twoheadleftarrow}{3}{arrows}{"23}
660 \DeclareMathSymbol{\leftleftarrows}{3}{arrows}{"71}
661 \DeclareMathSymbol{\rightrightarrows}{3}{arrows}{"73}
662 \DeclareMathSymbol{\upuparrows}{3}{arrows}{"72}
663 \DeclareMathSymbol{\downdownarrows}{3}{arrows}{"74}
664 \DeclareMathSymbol{\upharpoonright}{3}{arrows}{"75}
665 \DeclareMathSymbol{\downharpoonright}{3}{arrows}{"77}
666 \DeclareMathSymbol{\upharpoonleft}{3}{arrows}{"76}
667 \DeclareMathSymbol{\downharpoonleft}{3}{arrows}{"78}
668 \DeclareMathSymbol{\rightarrowtail}{3}{arrows}{"29}
669 \DeclareMathSymbol{\leftarrowtail}{3}{arrows}{"28}
670 \DeclareMathSymbol{\leftrightarrows}{3}{arrows}{"6E}
671 \DeclareMathSymbol{\rightleftarrows}{3}{arrows}{"6D}

672 \DeclareMathSymbol{\Lsh}{3}{arrows}{"7B}
673 \DeclareMathSymbol{\Rsh}{3}{arrows}{"7D}
674 \DeclareMathSymbol{\rightsquigarrow}{3}{arrows}{"8E}
675 \DeclareMathSymbol{\leftsquigarrow}{3}{arrows}{"8D}
676 \DeclareMathSymbol{\leftrightsquigarrow}{3}{arrows}{"91}
677 \DeclareMathSymbol{\looparrowleft}{3}{arrows}{"3F}
678 \DeclareMathSymbol{\looparrowright}{3}{arrows}{"40}
679 \DeclareMathSymbol{\circeq}{3}{symbols}{"D0}
680 \DeclareMathSymbol{\succsim}{3}{symbols}{"E1}
681 \DeclareMathSymbol{\gtrsim}{3}{symbols}{"DD}
682 \DeclareMathSymbol{\gtrapprox}{3}{letters}{"DB}
683 \DeclareMathSymbol{\multimap}{3}{letters}{"C7}
684 \DeclareMathSymbol{\image}{3}{letters}{"C6}
685 \DeclareMathSymbol{\original}{3}{letters}{"C5}
686 \DeclareMathSymbol{\therefore}{3}{symbols}{"90}
687 \DeclareMathSymbol{\because}{3}{symbols}{"91}
688 \DeclareMathSymbol{\doteqdot}{3}{symbols}{"CA}
689 \DeclareMathSymbol{\triangleq}{3}{symbols}{"D5}
690 \DeclareMathSymbol{\precsim}{3}{symbols}{"E0}
691 \DeclareMathSymbol{\lesssim}{3}{symbols}{"DC}
692 \DeclareMathSymbol{\lessapprox}{3}{letters}{"DA}
693 \DeclareMathSymbol{\eqslantless}{3}{letters}{"E2}
694 \DeclareMathSymbol{\eqslantgtr}{3}{letters}{"E3}
695 \DeclareMathSymbol{\curlyeqprec}{3}{letters}{"E6}
696 \DeclareMathSymbol{\curlyeqsucc}{3}{letters}{"E7}
697 \DeclareMathSymbol{\preccurlyeq}{3}{letters}{"E4}
698 \DeclareMathSymbol{\leqq}{3}{symbols}{"DA}
699 \DeclareMathSymbol{\leqslant}{3}{letters}{"E0}
700 \DeclareMathSymbol{\lessgtr}{3}{symbols}{"DE}
701 \DeclareMathSymbol{\backprime}{0}{letters}{"C8}
702 \DeclareMathSymbol{\axisshort}{0}{arrows}{"39}
703 \DeclareMathSymbol{\risingdotseq}{3}{symbols}{"CC}
704 \DeclareMathSymbol{\fallingdotseq}{3}{symbols}{"CB}
705 \DeclareMathSymbol{\succcurlyeq}{3}{letters}{"E5}
706 \DeclareMathSymbol{\geqq}{3}{symbols}{"DB}
707 \DeclareMathSymbol{\geqslant}{3}{letters}{"E1}
708 \DeclareMathSymbol{\gtrless}{3}{symbols}{"DF}
709 \let\sqsubset\undefined
710 \let\sqsupset\undefined
711 \DeclareMathSymbol{\sqsubset}{3}{symbols}{"E4}
712 \DeclareMathSymbol{\sqsupset}{3}{symbols}{"E5}
713 \DeclareMathSymbol{\vartriangleright}{3}{letters}{"2E}
714 \DeclareMathSymbol{\vartriangleleft}{3}{letters}{"2F}
715 \DeclareMathSymbol{\trianglerighteq}{3}{symbols}{"F5}
716 \DeclareMathSymbol{\trianglelefteq}{3}{symbols}{"F4}
717 \DeclareMathSymbol{\bigstar}{0}{arrows}{"AB}
718 \DeclareMathSymbol{\between}{3}{letters}{"F2}
719 \DeclareMathSymbol{\blacktriangledown}{0}{arrows}{"07}
720 \DeclareMathSymbol{\blacktriangleright}{3}{letters}{"F1}
721 \DeclareMathSymbol{\blacktriangleleft}{3}{letters}{"F0}

722 \DeclareMathSymbol{\arrowaxisright}{0}{arrows}{"37}
723 \DeclareMathSymbol{\arrowaxisleft}{0}{arrows}{"36}
724 \DeclareMathSymbol{\vartriangle}{3}{arrows}{"04}
725 \DeclareMathSymbol{\blacktriangle}{0}{arrows}{"05}
726 \DeclareMathSymbol{\triangledown}{0}{arrows}{"06}
727 \DeclareMathSymbol{\eqcirc}{3}{symbols}{"CF}
728 \DeclareMathSymbol{\lesseqgtr}{3}{letters}{"E8}
729 \DeclareMathSymbol{\gtreqless}{3}{letters}{"E9}
730 \DeclareMathSymbol{\lesseqqgtr}{3}{letters}{"EA}
731 \DeclareMathSymbol{\gtreqqless}{3}{letters}{"EB}
732 \DeclareMathSymbol{\Rrightarrow}{3}{arrows}{"6C}
733 \DeclareMathSymbol{\Lleftarrow}{3}{arrows}{"6A}
734 \DeclareMathSymbol{\veebar}{2}{letters}{"D2}
735 \DeclareMathSymbol{\barwedge}{2}{symbols}{"F6}
736 \DeclareMathSymbol{\angle}{0}{symbols}{"8B}
737 \DeclareMathSymbol{\measuredangle}{0}{symbols}{"8C}
738 \DeclareMathSymbol{\sphericalangle}{0}{symbols}{"8D}
739 \DeclareMathSymbol{\varpropto}{3}{symbols}{"2F} % ?
740 \DeclareMathSymbol{\smallsmile}{3}{letters}{"5E} % ?
741 \DeclareMathSymbol{\smallfrown}{3}{letters}{"5F} % ?
742 \DeclareMathSymbol{\Subset}{3}{symbols}{"F8}
743 \DeclareMathSymbol{\Supset}{3}{symbols}{"F9}
744 \DeclareMathSymbol{\Cup}{2}{symbols}{"FA}
745 \DeclareMathSymbol{\Cap}{2}{symbols}{"FB}
746 \DeclareMathSymbol{\curlywedge}{2}{symbols}{"84}
747 \DeclareMathSymbol{\curlyvee}{2}{symbols}{"85}
748 \DeclareMathSymbol{\leftthreetimes}{2}{letters}{"D0}
749 \DeclareMathSymbol{\rightthreetimes}{2}{letters}{"D1}
750 \DeclareMathSymbol{\subseteqq}{3}{letters}{"EE}
751 \DeclareMathSymbol{\supseteqq}{3}{letters}{"EF}
752 \DeclareMathSymbol{\bumpeq}{3}{symbols}{"C8}
753 \DeclareMathSymbol{\Bumpeq}{3}{symbols}{"C7}
754 \DeclareMathSymbol{\lll}{3}{letters}{"DE}
755 \DeclareMathSymbol{\ggg}{3}{letters}{"DF}
756 \DeclareMathSymbol{\circledS}{0}{letters}{"CA}
757 \DeclareMathSymbol{\pitchfork}{3}{letters}{"F3}
758 \DeclareMathSymbol{\dotplus}{2}{symbols}{"89}
759 \DeclareMathSymbol{\backsim}{3}{letters}{"F8}
760 \DeclareMathSymbol{\backsimeq}{3}{letters}{"F9}
761 \DeclareMathSymbol{\complement}{0}{letters}{"94}
762 \DeclareMathSymbol{\intercal}{2}{letters}{"D9}
763 \DeclareMathSymbol{\circledcirc}{2}{symbols}{"E6}
764 \DeclareMathSymbol{\circledast}{2}{symbols}{"E7}
765 \DeclareMathSymbol{\circleddash}{2}{letters}{"CC}

MSBM* equivalents

766 \DeclareMathSymbol{\lvertneqq}{3}{arrows}{"DE}
767 \DeclareMathSymbol{\gvertneqq}{3}{arrows}{"DF}
768 \DeclareMathSymbol{\nleq}{3}{arrows}{"9C}
769 \DeclareMathSymbol{\ngeq}{3}{arrows}{"9D}

770 \DeclareMathSymbol{\nless}{3}{arrows}{"9A}
771 \DeclareMathSymbol{\ngtr}{3}{arrows}{"9B}
772 \DeclareMathSymbol{\nprec}{3}{arrows}{"E5}
773 \DeclareMathSymbol{\nsucc}{3}{arrows}{"E6}
774 \DeclareMathSymbol{\lneqq}{3}{arrows}{"DC}
775 \DeclareMathSymbol{\gneqq}{3}{arrows}{"DD}
776 \DeclareMathSymbol{\nleqslant}{3}{arrows}{"D6}
777 \DeclareMathSymbol{\ngeqslant}{3}{arrows}{"D7}
778 \DeclareMathSymbol{\lneq}{3}{arrows}{"DA}
779 \DeclareMathSymbol{\gneq}{3}{arrows}{"DB}
780 \DeclareMathSymbol{\npreceq}{3}{arrows}{"E7}
781 \DeclareMathSymbol{\nsucceq}{3}{arrows}{"E8}
782 \DeclareMathSymbol{\precnsim}{3}{arrows}{"EB}
783 \DeclareMathSymbol{\succnsim}{3}{arrows}{"EC}
784 \DeclareMathSymbol{\lnsim}{3}{arrows}{"E0}
785 \DeclareMathSymbol{\gnsim}{3}{arrows}{"E2}
786 \DeclareMathSymbol{\nleqq}{3}{arrows}{"D8}
787 \DeclareMathSymbol{\ngeqq}{3}{arrows}{"D9}
788 \DeclareMathSymbol{\precneqq}{3}{arrows}{"E9}
789 \DeclareMathSymbol{\succneqq}{3}{arrows}{"EA}
790 \DeclareMathSymbol{\precnapprox}{3}{arrows}{"ED}
791 \DeclareMathSymbol{\succnapprox}{3}{arrows}{"EE}
792 \DeclareMathSymbol{\lnapprox}{3}{arrows}{"E3}
793 \DeclareMathSymbol{\gnapprox}{3}{arrows}{"E4}
794 \DeclareMathSymbol{\nsim}{3}{arrows}{"96}
795 \DeclareMathSymbol{\ncong}{3}{arrows}{"99}
796 \DeclareMathSymbol{\diagup}{3}{arrows}{"0B}
797 \DeclareMathSymbol{\diagdown}{3}{arrows}{"0C}
798 \DeclareMathSymbol{\varsubsetneq}{3}{arrows}{"D0}
799 \DeclareMathSymbol{\varsupsetneq}{3}{arrows}{"D1}
800 \DeclareMathSymbol{\nsubseteqq}{3}{arrows}{"CA}
801 \DeclareMathSymbol{\nsupseteqq}{3}{arrows}{"CB}
802 \DeclareMathSymbol{\subsetneqq}{3}{arrows}{"CE}
803 \DeclareMathSymbol{\supsetneqq}{3}{arrows}{"CF}
804 \DeclareMathSymbol{\varsubsetneqq}{3}{arrows}{"D2}
805 \DeclareMathSymbol{\varsupsetneqq}{3}{arrows}{"D3}
806 \DeclareMathSymbol{\subsetneq}{3}{arrows}{"CC}
807 \DeclareMathSymbol{\supsetneq}{3}{arrows}{"CD}
808 \DeclareMathSymbol{\nsubseteq}{3}{arrows}{"C8}
809 \DeclareMathSymbol{\nsupseteq}{3}{arrows}{"C9}
810 \DeclareMathSymbol{\nparallel}{3}{arrows}{"F7}
811 \DeclareMathSymbol{\nmid}{3}{arrows}{"F6}
812 \DeclareMathSymbol{\nshortmid}{3}{arrows}{"F4}
813 \DeclareMathSymbol{\nshortparallel}{3}{arrows}{"F5}
814 \DeclareMathSymbol{\nvdash}{3}{arrows}{"F8}
815 \DeclareMathSymbol{\nVdash}{3}{arrows}{"FA}
816 \DeclareMathSymbol{\nvDash}{3}{arrows}{"F9}
817 \DeclareMathSymbol{\nVDash}{3}{arrows}{"FB}
818 \DeclareMathSymbol{\ntrianglerighteq}{3}{arrows}{"F2}
819 \DeclareMathSymbol{\ntrianglelefteq}{3}{arrows}{"F1}

820 \DeclareMathSymbol{\ntriangleleft}{3}{arrows}{"EF}
821 \DeclareMathSymbol{\ntriangleright}{3}{arrows}{"F0}
822 \DeclareMathSymbol{\nleftarrow}{3}{arrows}{"32}
823 \DeclareMathSymbol{\nrightarrow}{3}{arrows}{"33}
824 \DeclareMathSymbol{\nLeftarrow}{3}{arrows}{"66}
825 \DeclareMathSymbol{\nRightarrow}{3}{arrows}{"68}
826 \DeclareMathSymbol{\nLeftrightarrow}{3}{arrows}{"67}
827 \DeclareMathSymbol{\nleftrightarrow}{3}{arrows}{"34}
828 \DeclareMathSymbol{\divideontimes}{2}{letters}{"F7}
829 \DeclareMathSymbol{\varnothing}{0}{letters}{"9C}
830 \DeclareMathSymbol{\nexists}{0}{arrows}{"20}
831 \DeclareMathSymbol{\Finv}{0}{letters}{"90}
832 \DeclareMathSymbol{\Game}{0}{letters}{"91}
833 \let\mho\undefined
834 \DeclareMathSymbol{\mho}{0}{letters}{"92}
835 \DeclareMathSymbol{\simeq}{3}{symbols}{"27}
836 \DeclareMathSymbol{\eqsim}{3}{symbols}{"99}
837 \DeclareMathSymbol{\beth}{0}{letters}{"95}
838 \DeclareMathSymbol{\gimel}{0}{letters}{"96}
839 \DeclareMathSymbol{\daleth}{0}{letters}{"97}
840 \DeclareMathSymbol{\lessdot}{3}{letters}{"DC}
841 \DeclareMathSymbol{\gtrdot}{3}{letters}{"DD}
842 \DeclareMathSymbol{\ltimes}{2}{letters}{"CE}
843 \DeclareMathSymbol{\rtimes}{2}{letters}{"CF}
844 \DeclareMathSymbol{\shortmid}{3}{letters}{"F4}
845 \DeclareMathSymbol{\shortparallel}{3}{letters}{"F5}
846 \DeclareMathSymbol{\smallsetminus}{2}{letters}{"D8} %?
847 \DeclareMathSymbol{\thicksim}{3}{symbols}{"18} %?
848 \DeclareMathSymbol{\thickapprox}{3}{symbols}{"19} %?
849 \DeclareMathSymbol{\approxeq}{3}{symbols}{"9D}
850 \DeclareMathSymbol{\succapprox}{3}{letters}{"ED}
851 \DeclareMathSymbol{\precapprox}{3}{letters}{"EC}
852 \DeclareMathSymbol{\curvearrowleft}{3}{arrows}{"87}
853 \DeclareMathSymbol{\curvearrowright}{3}{arrows}{"88}
854 \DeclareMathSymbol{\digamma}{0}{letters}{"46} %?
855 \DeclareMathSymbol{\varkappa}{0}{letters}{"9B}
856 \DeclareMathSymbol{\Bbbk}{0}{arrows}{"6B}
857 \DeclareMathSymbol{\hslash}{0}{letters}{"9D}
858 \DeclareMathSymbol{\hbar}{0}{arrows}{"1B}
859 \DeclareMathSymbol{\backepsilon}{3}{letters}{"FB} %?
860 \DeclareMathSymbol{\dashrightarrow}{0}{arrows}{"3A}
861 \DeclareMathSymbol{\dashleftarrow}{0}{arrows}{"38}
862 \DeclareMathSymbol{\dashuparrow}{0}{arrows}{"39}
863 \DeclareMathSymbol{\dashdownarrow}{0}{arrows}{"3B}
864 \DeclareMathDelimiter\ulcorner{4}{arrows}{"70}{arrows}{"70}
865 \DeclareMathDelimiter\urcorner{5}{arrows}{"71}{arrows}{"71}
866 \DeclareMathDelimiter\llcorner{4}{arrows}{"78}{arrows}{"78}
867 \DeclareMathDelimiter\lrcorner{5}{arrows}{"79}{arrows}{"79}
868 \edef\checkmark{\noexpand\mathhexbox{\hexnumber@\symarrows}AC}
869 \edef\circledR{\noexpand\mathhexbox{\hexnumber@\symletters}C9}

870 \edef\maltese{\noexpand\mathhexbox{\hexnumber@\symletters}CB}

Changes to default for Leftrightarrow. I (SPQR) don’t like 22C, so:

871 \let\Leftrightarrow\undefined
872 \DeclareMathSymbol{\Leftrightarrow}{3}{arrows}{"61}

Override AMS logo, just to ensure we don’t use any CM fonts!

873 \def\AmS{{\protect\AmSfont
874 A\kern-.1667em\lower.5ex\hbox{M}\kern-.125emS}}

875 ⟨lucidabright | lucbmath⟩\def\AmSfont{\usefont{OMS}{\ifnum\Lucida@names=0 hlcy\else lby \fi}{m}{n}} 876 ⟨/lucbmath⟩ 877 ⟨∗lucbrb⟩

878 \ProvidesPackage{lucidbrb}[\filedate\space\fileversion\space
879 Lucida Bright (Berry names) PSNFSS2e package]
880 \DeclareOption*{\PassOptionsToPackage{\CurrentOption}{lucbr}}
881 \ProcessOptions
882 \RequirePackage{lucbr}

883 ⟨/lucbrb⟩ 884 ⟨∗lucbry⟩

885 \ProvidesPackage{lucidbry}[\filedate\space\fileversion\space
886 Lucida Bright (Y \& Y names) PSNFSS2e package]
887 \DeclareOption*{\PassOptionsToPackage{\CurrentOption}{lucbr}}
888 \ProcessOptions
889 \RequirePackage[yy]{lucbr}

890 ⟨/lucbry⟩

A test file for the Lucida fonts.

891 ⟨∗lucfont⟩

892 %    \begin{macrocode}
893 \documentclass{article}
894 \usepackage{t1enc}
895 \begin{document}
896 \title{All the Lucida text fonts}
897 \author{prepared by Sebastian Rahtz}
898 \date{February 19th 1995}
899 \maketitle
900 \def\test#1#2#3#4#5{%
901 \item[#1/#2/#3]#4 (#5):
902 {\fontfamily{#1}\fontseries{#2}\fontshape{#3}\selectfont
903 Animadversion for a giraffe costs \pounds123. Wa\ss\ ist
904 das f\"ur ein Klopf?
905 We are often na{\"\i}ve vis-\`{a}-vis
906 the d{\ae}monic ph{\oe}nix's official r\^{o}le in fluffy souffl\'{e}s}

907 }

908 \begin{description}
909 \test{hlx}{db}{it}{hlxdi8t}{LucidaFax-DemiItalic}
910 \test{hlx}{db}{n}{hlxd8t}{LucidaFax-Demi}
911 \test{hlx}{m}{it}{hlxrir8t}{LucidaFax-Italic}
912 \test{hlx}{m}{n}{hlxr8t}{LucidaFax}
913 \test{hlh}{db}{it}{hlcdib8t}{LucidaBright-DemiItalic}
914 \test{hlh}{db}{n}{hlcdb8t}{LucidaBright-Demi}

915 \test{hlh}{m}{it}{hlcrib8t}{LucidaBright-Italic}
916 \test{hlh}{m}{n}{hlcrb8t}{LucidaBright}
917 \test{hlce}{m}{it}{hlcrie8t}{LucidaCalligraphy-Italic}
918 \test{hlcf}{m}{n}{hlcrf8t}{LucidaBlackletter}
919 \test{hlcn}{m}{it}{hlcrin8t}{LucidaCasual-Italic}
920 \test{hlcn}{m}{n}{hlcrn8t}{LucidaCasual}
921 \test{hlst}{b}{n}{hlsbt8t}{LucidaSans-TypewriterBold}
922 \test{hlst}{b}{sl}{hlsbot8t}{LucidaSans-TypewriterBoldOblique}
923 \test{hls}{b}{it}{hlsbi8t}{LucidaSans-BoldItalic}
924 \test{hls}{b}{n}{hlsb8t}{LucidaSans-Bold}
925 \test{hls}{db}{it}{hlsdi8t}{LucidaSans-DemiItalic}
926 \test{hls}{db}{n}{hlsd8t}{LucidaSans-Demi}
927 \test{hls}{m}{it}{hlsri8t}{LucidaSans-Italic}
928 \test{hls}{m}{n}{hlsr8t}{LucidaSans}
929 \test{hlct}{b}{n}{hlcbt8t}{LucidaTypewriterBold}
930 \test{hlct}{b}{sl}{hlcbot8t}{LucidaTypewriterOblique}
931 \test{hlcw}{m}{it}{hlcriw8t}{LucidaHandwriting-Italic}
932 \end{description}
933 \end{document}

934 ⟨/lucfont⟩